### v1

# BOOST: Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique

Abstract
The performance of Bayesian optimization (BO), a highly sample-efficient method for expensive black-box problems, is critically governed by the selection of its hyperparameters, including the kernel and acquisition functions. This presents a significant practical challenge: an inappropriate combination of these can lead to poor performance and wasted evaluations. While individual improvements to kernel functions (e.g., tree-based kernels, deep kernel learning) and acquisition functions (e.g., multi-step lookahead, tree-based planning) have been actively explored, the joint and autonomous selection of the best pair of these fundamental hyperparameters has been largely overlooked. This forced practitioners to rely on heuristics or costly manual training. In this work, we propose a simple yet effective framework, BOOST (Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique), that automates this selection. BOOST utilizes a lightweight, offline evaluation stage to predict the performance of various kernel-acquisition function pairs and identify the most suitable configuration before committing to the expensive evaluation process. BOOST partitions data-in-hand into two subsets: a reference subset and a query subset, and it prepares all possible kernel-acquisition pairs from the user’s chosen candidates. For each configuration, BOOST conducts internal BO runs using the reference subset, evaluating how effectively each pair guides the search toward the optimum in the unknown query subset, thereby identifying the configuration with the best retrospective performance for future optimization. Experiments on both synthetic benchmark functions and real-world hyperparameter optimization tasks demonstrate that BOOST consistently outperforms standard BO approaches with fixed hyperparameters, highlighting its effectiveness and robustness in diverse problem landscapes.
Introduction
Black-box optimization is a common challenge in scientific and engineering problems, especially when each function evaluation is time-consuming or costly (Jones et al. 1998). In this context, a black box refers to situations where the objective function cannot be modeled and only input-output relationships are observable, with no access to gradients or knowledge of the underlying system. These challenges necessitate efficient methods that minimize the number of evaluations to approach optimal solutions.
The fundamental challenge in black-box optimization lies in managing uncertainty about the objective function’s behavior in unexplored regions. The high cost of function evaluations means that only a limited number of observations can be obtained, requiring optimization strategies to make informed decisions under uncertainty with sparse data. Bayesian Optimization (BO) addresses this challenge by employing a principled probabilistic approach that combines a surrogate model—most commonly a Gaussian Process (GP) (Williams and Rasmussen 2006)—with an acquisition function to guide the search process efficiently (Frazier 2018; Shahriari et al. 2015). GP provides uncertainty estimates of the objective function, while the acquisition function leverages these estimates to determine the next promising point to evaluate. This combination allows BO to systematically balance exploration and exploitation, leading to data-efficient optimization (Snoek et al. 2012).
Given these advantages, BO is broadly used for costly experiments, including material synthesis (Pruksawan et al. 2019; Shields et al. 2021), optimizing process variables (Greenhill et al. 2020; Park et al. 2024a, b), and machine learning hyperparameter optimizations (Chen et al. 2018; Falkner et al. 2018; Feurer et al. 2015; Kandasamy et al. 2018; Shen et al. 2023; Snoek et al. 2012; White et al. 2021).
Despite its widespread application, the performance of BO critically depends on the proper choice of two hyperparameters: the choice of kernel function used in the GP and the choice of acquisition function guiding the subsequent sample point (Ginsbourger et al. 2008; Hoffman et al. 2011; Roman et al. 2019; Snoek et al. 2012; Vasconcelos et al. 2019). Nevertheless, these hyperparameters are typically fixed throughout the optimization process, generally chosen arbitrarily or based on prior domain knowledge (Genton 2001; Wilson and Adams 2013), with little attention paid to their selection or adaptation (Roman et al. 2019). This oversight neglects an important insight: the optimal combination of kernel and acquisition function can vary significantly depending on the structure and complexity of the objective function. However, the black-box nature of these problems creates a fundamental paradox: practitioners must select hyperparameters that match the function’s characteristics without knowing those characteristics beforehand.
Some existing approaches have attempted to address this gap by adaptive selection of kernel functions (Ginsbourger et al. 2008; Roman et al. 2019) or acquisition functions (Hoffman et al. 2011; Vasconcelos et al. 2019). However, these methods typically rely on local information, such as the outcome of a single query point or uncertainty estimates from the surrogate model, which may not always reflect the true structure of the objective function, particularly in noisy or complex optimization landscapes. Another active research direction is individual development of kernel functions (Wilson et al. 2016; Boyne et al. 2025) or acquisition functions (Wu and Frazier 2019; Lam et al. 2016; Song et al. 2022; Cheon et al. 2024). These methods, however, focus on improving either the kernel or the acquisition function while leaving the other fixed, without considering their joint selection.
To address these challenges, we propose BOOST (Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique), a simple yet effective method that identifies the most suitable kernel–acquisition pair based entirely on the previously evaluated points. By performing offline evaluations using the data-in-hand, BOOST avoids dependence on uncertain model estimates and local improvements.
For each BO iteration, BOOST selects the suitable kernel-acquisition function pair through a three-stage process: data partitioning, performance evaluation, and recommendation. We first partition the data-in-hand into two subsets: one used as the reference set to construct a GP for the internal BO process, and the other treated as a query set which serves as unexplored regions of internal BO where new query points are generated.
For each candidate kernel-acquisition pair predefined by users, BOOST fits a GP to the reference subset and selects the next query point in the query subset. This internal BO is repeated until it reaches the target value in the query subset. The number of iterations required by each pair to reach the target is recorded, and the pair with minimum iteration number is selected for the actual BO iteration.
This approach breaks the intrinsic paradox of BO, selecting appropriate hyperparameters without prior knowledge of the objective function, by leveraging the full structure embedded in the available data, rather than relying on noisy model uncertainty. BOOST thus converts available data into actionable insights, enabling robust and informed hyperparameter selection.
Experiments on both synthetic benchmark functions and real-world hyperparameter optimization tasks show that BOOST consistently outperforms fixed-hyperparameter baselines, highlighting the importance of careful BO hyperparameter selection for achieving significant performance gains.
In summary, the key contributions of BOOST are as follows:
  • Systematic approach for joint kernel–acquisition function selection.
  • Breaking the black-box paradox without prior knowledge.
Preliminaries
Assume we aim to find the global maximum,
$x^{*}=\arg\max_{x\in\mathcal{X}}f(x),$
where $f$ is an expensive black-box function. In this setting, we have no access to gradients or structural information about $f$. The only available information is a set of observed input–output pairs, known as the data-in-hand, $D_{n}=\{x_{i},f(x_{i})\}_{i=1}^{n}$.
Bayesian Optimization (BO) addresses this challenge by leveraging a probabilistic surrogate model. BO proceeds iteratively in two main steps. First, a Gaussian Process (GP) is constructed based on data-in-hand, using a kernel function that models the relationship between data-in-hand and unobserved points in the search space. The GP serves as a surrogate model for the unknown objective function. Second, an acquisition function, which leverages the GP’s predictive distribution, is utilized to determine the next most promising point to evaluate. This process is repeated until a stopping criterion is met, such as reaching a target performance or evaluation budget. This framework enables efficient optimization by balancing the exploration of uncertain regions with the exploitation of promising areas.
The performance of Bayesian Optimization largely depends on two key components: the kernel function and the acquisition function. The kernel function governs how the GP makes predictions by defining the correlation structure between input points. It controls the smoothness, scale, and other structural assumptions embedded in the surrogate model, thereby directly influencing the accuracy of predictions and the model’s uncertainty estimates.
The acquisition function, by contrast, determines the optimization strategy by quantifying the utility of candidate points based on the GP’s predictions. It guides the search process by balancing the trade-off between exploring uncertain regions and exploiting promising areas likely to yield high-performing outcomes. Therefore, selecting an appropriate kernel and acquisition function is crucial to the success of Bayesian Optimization.
Gaussian Process (GP)
GP is a statistical model used to predict the behavior of an unknown objective function $f$ on domain $\mathcal{X}$, based on given dataset $D_{n}=\{x_{i},f(x_{i})\}_{i=1}^{n}$.
GP defines a distribution over functions, characterized by its mean function $m:\mathcal{X}\rightarrow\mathbf{R}$ and kernel function (or covariance function) $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbf{R}$:
$f(x)\sim\mathcal{GP}(m(x),k(x,x^{\prime}))$ (1)
The prediction of a given point $x^{\prime}\in\mathcal{X}$ is
$f(x^{\prime}|D_{n})\sim\mathcal{N}(\mu(x^{\prime}),\sigma^{2}(x^{\prime}))$ (2)
where
$\mu(x^{\prime})=k(x^{\prime})^{\top}K^{-1}(f(X)-m(X))+m(x^{\prime})$ (3)
$\sigma^{2}(x^{\prime})=k(x^{\prime},x^{\prime})-k(x^{\prime})^{\top}K^{-1}k(x^{\prime})$ (4)
Here, $k(x^{\prime})\in\mathbf{R}^{n}$ denotes the covariance vector between $x^{\prime}$ and all $x_{i}$ in the dataset, and $K$ denotes the covariance matrix between all pairs of $x_{i}$, i.e., $K_{ij}=k(x_{i},x_{j})$. Also, $f(X)=[f(x_{1}),f(x_{2}),\dots,f(x_{n})]^{\top}$ denotes the vector of observed function values.
Note that the kernel function determines the GP’s modeling capability by encoding assumptions about the function’s structure. The following are widely used kernel functions.
$k_{\mathrm{Mat\acute{e}rn}}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,r}{\ell}\right)^{\nu}K_{\nu}\!\left(\frac{\sqrt{2\nu}\,r}{\ell}\right)$ (5)
$k_{\mathrm{RBF}}(r)=\exp\left(-\frac{r^{2}}{2l^{2}}\right)$ (6)
$k_{\mathrm{RQ}}(r)=\left(1+\frac{r^{2}}{2\alpha l^{2}}\right)^{-\alpha}$ (7)
where $r=\|x-x^{\prime}\|$ is the Euclidean distance between two input points, $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind.
The Radial Basis Function (RBF) kernel (Lowe and Broomhead 1988) assumes an infinitely smooth function. In contrast, the Matérn kernel (Matérn 1960; Williams and Rasmussen 2006) offers flexibility in modeling less smooth functions, with $\nu=$ 3/2 and 5/2 being widely used for real-world tasks. The Rational Quadratic (RQ) kernel (Williams and Rasmussen 2006) is capable of modeling functions with multiple characteristic length scales, with $\alpha=2$ being widely used. Throughout this paper, we fix the parameter $\alpha$ of the RQ kernel to 2. Given this variability, the choice of kernel can have a substantial impact on BO performance depending on the structure of the objective function.
Acquisition function
Based on the GP’s predictions, the acquisition function balances exploration and exploitation to suggest the next query point,
$x_{n+1}\leftarrow\arg\max_{x\in\mathcal{X}}\alpha(x).$
Among various acquisition functions, Expected Improvement (EI) (Jones et al. 1998) is one of the most widely used. It computes the expected improvement over the current best observation, thereby naturally balancing exploration (through predictive uncertainty) and exploitation (through mean estimate). The following are widely used acquisition functions.
$\alpha_{\mathrm{EI}}(x)=\mathbf{E}\left[\max(f(x)-f^{+},0)\right]$ (8)
$\alpha_{\mathrm{PI}}(x)=\mathbf{P}(f(x)>f^{+})$ (9)
$\alpha_{\mathrm{UCB}}(x)=\mu(x)+\beta\sigma(x)$ (10)
$\alpha_{\mathrm{PM}}(x)=\mu(x)$ (11)
where $f^{+}$ denotes the best observed function value so far.
Different acquisition functions correspond to different search strategies. Probability of Improvement (PI) (Kushner 1964) focuses solely on the probability of achieving better outcomes than the current best, often leading to overly greedy behavior if uncertainty is underestimated. Upper Confidence Bound (UCB) (Srinivas et al. 2010) adds a weighted uncertainty term to the predicted mean, enabling more principled exploration. The parameter $\beta$ controls the exploration–exploitation trade-off: a low $\beta$ (e.g., 0.1) encourages exploitation behavior, often more than PI, while a high $\beta$ (e.g., 5) encourages aggressive exploration, surpassing even EI in exploratory behavior (Papenmeier et al. 2025). Since no standard value exists for $\beta$, we adopt $\beta=0.1$ with reference to the values used in some recent studies (Dogan and Prestwich 2022; Tian et al. 2024). In contrast, Posterior Mean (PM) chooses the next point solely based on the GP’s predicted mean value. While this may work well when the model is highly confident, it completely ignores uncertainty and thus represents the most greedy strategy among common acquisition functions.
Related Works
Adaptive Kernel and Acquisition Selection in BO
Various approaches have been proposed to address the challenge of selecting appropriate kernel and acquisition functions in Bayesian Optimization (BO), recognizing their strong influence on performance. GP-Hedge (Hoffman et al. 2011) addresses the acquisition selection problem by treating it as a multi-armed bandit, adaptively adjusting the probability distribution over acquisition functions based on the surrogate model’s predicted value at each candidate point they propose. However, this approach heavily relies on uncertain, model-driven improvements from single points, leading to potential instability, especially in the early stages of optimization. No-PASt-BO (Vasconcelos et al. 2019) mitigates this by incorporating historical data with a memory factor, yet still suffers from reliance on uncertain predictive values and does not address kernel adaptation.
Similarly, adaptive kernel selection approaches have been explored. Ginsbourger et al. (2008) proposed using discrete mixtures of kernels within Gaussian Process-based optimization (originally referred to as Kriging), where the surrogate model is constructed from a weighted combination of multiple kernels. Roman et al. (2019), on the other hand, developed a strategy to select the best-performing kernel among multiple candidates by evaluating them in parallel. While these approaches reduce kernel misspecification risks, they both rely on local surrogate information for decision-making and neglect the influence of acquisition function choice, which can limit their overall robustness and adaptability in complex optimization settings.
In contrast to these approaches that focus on individual components, some recent work has recognized the need for joint selection. The fundamental challenge in black-box optimization is that practitioners must choose appropriate kernel and acquisition functions that match the function’s characteristics without knowing those characteristics beforehand, making simultaneous optimization of both components particularly difficult. To address this issue, Xue et al. (2016) proposed a data-based strategy that utilizes the data-in-hand to jointly select the most appropriate surrogate model and acquisition function for material design tasks. While this approach represents a significant step toward breaking the conventional paradox of selecting hyperparameters without prior knowledge—by transforming available data into actionable insights—their method selects hyperparameters only at the beginning and does not update them during subsequent iterations. This static approach fails to leverage the additional information that becomes available as the optimization progresses, limiting its adaptability to the evolving understanding of the objective function.
Building upon this foundation, BOOST introduces an adaptive data-based strategy that utilizes all previously evaluated points to retrospectively assess candidate configurations at each iteration. Unlike previous adaptive kernel function or adaptive acquisition function studies, BOOST provides several key advantages: it requires no additional function evaluations, pretraining, or assumptions about model uncertainty for joint hyperparameter selection. Moreover, by updating hyperparameter choices at every iteration using the continuously expanding data-in-hand, BOOST successfully addresses the conventional paradox of selecting kernel and acquisition functions in black-box settings resulting in superior performance.
Methods
This section outlines the structure of BOOST and how it leverages observed data to guide BO hyperparameter selection. We describe the overall procedure, followed by the details of each component. Without loss of generality, BOOST assumes minimization problems throughout this section.
Overall Process of BOOST
Figure 1: An Overview of BOOST architecture.
BOOST performs offline evaluation using data-in-hand to identify the most suitable kernel-acquisition pair from candidates predefined by practitioners. BOOST operates in three main steps:
  1. 1.
    Data Partitioning: The data-in-hand, $D_{n}$, is partitioned into two disjoint subsets: a reference subset, $r_{n}$, to construct the Gaussian Process (GP) for the internal BO process, and a query set, and $q_{n}$ which serves as unexplored regions of internel BO where new query points are generated.
  2. 2.
    Performance Evaluation: The optimization performance of each candidate kernel–acquisition pair is evaluated based on how quickly it discovers the target value $y_{target}$ within $q_{n}$, starting from $r_{n}$.
  3. 3.
    Recommendation: The kernel–acquisition pair that achieves the fastest convergence (i.e., reaches the target value in the fewest steps) is selected for the actual BO process.
The overview of BOOST is shown in Figure 1 and Algorithm 1.
Algorithm 1 Overall algorithm for BOOST
Input: The set of kernel function candidates $\mathcal{K}$, The set of acquisition function candidates $\mathcal{A}$, Data-in-hand $D_{n}$
Parameter: Constraint of initial size $(n_{\min},n_{\max})$, ratio $r$, Maximum iterations $T_{\max}$, Stopping criterion percentile $a$
Output: Best Kernel-Acquitision function pair ($k^{*}$, $\alpha^{*}$)
[…]
5:for all $(k_{i},\alpha_{j})$ in $\mathcal{K}\times\mathcal{A}$ do
6:  $R_{0}\leftarrow r_{n}=\{(x_{i},f(x_{i}))\}_{i=1}^{n_{init}}$
7:  $Q_{0}\leftarrow\{x\mid(x,f(x))\in q_{n}\}$
8:  $t\leftarrow 0$
9:  while $t<T_{\max}$ do
10:   Train a Gaussian Process using $k_{i}$ over $R_{t}$
11:   $x_{t+1}\leftarrow\arg\max_{x\in Q_{t}}\alpha_{j}(x)$
12:   $R_{t+1}\leftarrow R_{t}\cup\{(x_{t+1},f(x_{t+1}))\}$
13:   $Q_{t+1}\leftarrow Q_{t}\setminus\{x_{t+1}\}$
14:   $t\leftarrow t+1$
15:   if $f(x_{t+1})\leq y_{target}$ then
16:    break
17:   end if
18:  end while
19:  $t_{k_{i},\alpha_{j}}\leftarrow t$
20:end for
21:return $(k^{*},\alpha^{*})=\arg\min_{(k_{i},\alpha_{j})}t_{k_{i},\alpha_{j}}$
Data Partitioning
The data-in-hand, $D_{n}$, is partitioned into two disjoint subsets: $r_{n}$ and $q_{n}$. In BOOST, $r_{n}$, which is the reference subset for internal BO process, corresponds to the initial dataset used in the actual BO loop and is utilized to construct the GP at the beginning of the performance evaluation step.
To ensure that $r_{n}$ consists of informative and diverse samples while excluding points already better than the target value (or stopping criteria), $y_{target}$, we apply K-means clustering on $D_{n}\setminus\{(x,f(x))\mid f(x)\leq y_{target}\}$ to select representative samples that provide diverse coverage of the promising regions. From each cluster, the data point closest to the cluster center is selected, providing both representativeness and coverage of promising regions. In this study, we define the target value $y_{target}$ for stopping as the top 5th percentile of function values in $D_{n}$, while the size of $r_{n}$ is determined as $\lfloor|D_{n}|/3\rfloor$ with the constraint of $3\leq|r_{n}|\leq 20$ to balance the size of $r_{n}$ and $q_{n}$, where $q_{n}=D_{n}\setminus r_{n}$. The performance impact of different stopping criteria and data partitioning methods is analyzed in the Appendix.
All function values in $q_{n}$, $\{f(x)\mid x\in q_{n}\}$, are treated as undiscovered data to assess performance, mimicking the unexplored regions typically encountered in realistic black-box optimization scenarios.
Performance Evaluation
BOOST evaluates combinations from predefined candidate sets:
  • Kernel functions: {Matérn 3/2, Matérn 5/2, RBF, RQ}
  • Acquisition functions: {EI, PI, UCB ($\beta=0.1$), PM}
The mathematical definitions of these functions are provided in the Preliminaries section.
For each kernel–acquisition pair $(k_{i},\alpha_{j})$, BOOST performs the following steps:
  1. 1.
    Train a Gaussian Process using kernel $k_{i}$ over the reference subset $r_{n}$.
  2. 2.
    Using predicted mean value and uncertainty from GP in step 1, select next query point $x_{n+1}$ from $q_{n}$ that maximizes acquisition function $\alpha_{j}$, and move $(x_{n+1},f(x_{n+1}))$ from $q_{n}$ to $r_{n}$.
  3. 3.
    Repeat internal BO (i.e., step 1 and step 2) until:
    • The target value $y_{target}$ is reached, i.e., $f(x_{n+1})\leq y_{target}$, or
    • A maximum of 20 iterations is reached.
  4. 4.
    Record the number of iterations $t_{k_{i},\alpha_{j}}$ required to achieve the target.
  5. 5.
    Repeat steps 1–4 for all other kernel–acquisition combinations.
The internal BO process is limited to 20 iterations for computational efficiency. This efficiency is further enhanced by the fact that the data-based evaluation in BOOST is fully parallelizable, as each kernel–acquisition pair is assessed independently.
Recommendation
After evaluating all candidate combinations, BOOST selects the pair that required the fewest iterations to reach the target value:
$(k^{*},\alpha^{*})=\arg\min_{(k_{i},\alpha_{j})}t_{k_{i},\alpha_{j}}$
In cases where multiple combinations achieve the same minimum number of iterations, particularly when $D_{n}$ is small, BOOST breaks the tie using a predefined priority order:
  • Acquisition functions: EI $>$ PI $>$ UCB $>$ PM
  • Kernel functions: Matérn 3/2 $>$ Matérn 5/2 $>$ RBF $>$ RQ
This priority is based on the degree of exploration each acquisition function promotes. When data is scarce, exploration is essential to acquire informative samples that help reveal the structural characteristics of the objective function. This, in turn, enables the internal BO process to more effectively evaluate and distinguish between different kernel–acquisition combinations. Accordingly, we prioritize acquisition functions with stronger exploratory tendencies, ranking them as EI $>$ PI $>$ UCB $>$ PM. In particular, UCB is placed below PI because we adopt a low exploration coefficient $(\beta=0.1)$, which makes it behave almost greedily (Papenmeier et al. 2025; Dogan and Prestwich 2022; Tian et al. 2024).
Similarly, we prioritize Matérn kernels over RBF due to their greater flexibility in modeling non-smooth or complex objective functions, as they are recommended for many realistic problems where strong smoothness assumptions are unrealistic (Snoek et al. 2012; Williams and Rasmussen 2006; Stein 1999). We place RBF over RQ since RBF is the most widely used kernel, while RQ is a special case of kernels which can be written as a mixture of RBF kernels with various length scales (Williams and Rasmussen 2006).
The tie-breaking rule incorporates domain-inspired heuristics. However, the core selection mechanism in BOOST is data-driven, based on internal BO performance evaluation. The heuristic component is only applied after quantitative evaluation, and only when multiple options are equally optimal. This minimal use of heuristics helps guide the decision in underdetermined cases without compromising the overall principled framework of BOOST.
Figure 2: Optimization performance of BOOST and 16 fixed-kernel-acquisition methods on synthetic benchmark functions and HPO-B datasets. HPO-B tasks include Decision Trees (6D, space IDs: 5636, 5859), Random Forests (9D, space IDs: 7607, 7609), and XGBoost (16D, space ID: 5906). Each experiment uses 10 initial points followed by 90 optimization iterations (100 total evaluations).
Function Ackley Levy Rosenbrock SumSquares 5636 5859 7607 7609 5906 Avg. Rank
BOOST 1 2 1 2 2 1 2 3 1 1.67
Matérn32_EI 10 1 3 4 4 2 4 6 9 4.78
Matérn52_EI 5 4 2 3 11 12 7 5 4 5.89
Matérn52_PI 2 10 8 9 5 5 3 2 16 6.67
RQ_EI 12 3 7 1 3 15 6 8 8 7.00
Matérn32_PI 3 13 6 16 10 9 9 4 10 8.89
RBF_PI 17 7 17 5 15 17 1 1 2 9.11
RQ_PI 9 6 14 10 14 3 8 14 5 9.22
RBF_EI 16 5 12 7 1 16 5 7 17 9.56
RBF_UCB 13 8 11 6 13 11 12 15 6 10.56
RQ_PM 8 11 15 13 8 6 10 9 15 10.56
Matérn32_PM 14 16 10 17 7 8 13 10 3 10.89
Matérn32_UCB 15 14 9 15 6 4 15 16 7 11.22
RQ_UCB 7 12 13 12 9 14 11 13 11 11.33
Matérn52_UCB 4 15 4 11 16 13 16 12 12 11.44
Matérn52_PM 6 17 5 14 12 10 17 11 13 11.67
RBF_PM 11 9 16 8 17 7 14 17 14 12.56
Table 1: Final performance rankings (lower is better) based on regret after 100 samples across four synthetic functions and five HPO-B benchmarks. Numbers 5636–5906 indicate the search space IDs used in the HPO-B dataset. To ensure fair ranking, if the regrets after $n$ evaluations were tied, we compared regrets at $n-1$ to resolve ties.
Experiments
We conduct comprehensive experiments to evaluate BOOST’s performance across various discrete optimization scenarios. All benchmark tasks are designed in fully discretized spaces to align with practical problem settings and to demonstrate BOOST’s real-world applicability (González-Duque et al. 2024). We compare BOOST against fixed kernel–acquisition function combinations on both synthetic benchmark functions and real-world hyperparameter optimization tasks.
All optimization problems in this section are minimization problems. All experiments are repeated 10 times with different set of initial points to ensure statistical reliability.
Synthetic Benchmark Functions
We evaluate BOOST on four widely used 4-dimensional discrete synthetic benchmark functions: Ackley, Levy, Rosenbrock, and SumSquares (Surjanovic and Bingham 2025). Detailed discretization parameters are provided in the Appendix. We use Latin Hypercube Sampling (LHS) to select 10 initial points, ensuring well-distributed coverage of the search space.
Real-world machine learning hyperparameter optimization tasks
To assess the effectiveness of BOOST in real-world problems, we conduct experiments on five hyperparameter optimization tasks selected from the HPO-B dataset (Arango et al. 2021). These tasks involve Decision Trees (6D), Random Forests (9D), and XGBoost (16D) with varying search space dimensionalities. For each task, we randomly select 10 initial points from the available dataset, as the search space is predefined and sampling methods such as LHS are not applicable. Detailed experimental configurations and data preprocessing steps are provided in the Appendix.
Experimental Setup
We compare BOOST against 16 deterministic baseline methods, representing all possible combinations of the four kernel functions (Matérn 3/2, Matérn 5/2, RBF, RQ) and four acquisition functions (EI, PI, UCB, PM) used in our candidate set.
Results
As shown in Figure 2, BOOST consistently demonstrates competitive performance across both synthetic benchmark functions and real-world machine learning hyperparameter optimization tasks. Compared to other fixed hyperparameter methods, BOOST ranked within the top 3 across all test cases when comparing regret after 100 evaluations (10 initial samples followed by 90 BO iterations), highlighting its adaptability and effectiveness in diverse optimization scenarios. To ensure fair ranking, if the regrets after $n$ evaluations were tied, we compared regrets at $n-1$ to resolve ties.
BOOST achieved the best overall performance with an average rank of 1.67 out of 18 candidates (Table 1). The second-best method, Matérn 3/2 with EI, had an average rank of 4.78. This suggests that BOOST is not only competitive on simple problems but also robust across both simple and complex functions, consistently identifying suitable kernel and acquisition combinations for any possible search space.
Conclusion
In this study, we propose BOOST, a robust framework for selecting appropriate hyperparameters in Bayesian Optimization (BO). Motivated by the insight that the data-in-hand can help us to understand the shape and complexity of the objective function, BOOST performs internal BO loop on the available data to identify the kernel and acquisition function combination that is best suited to the target problem.
Unlike fixed hyperparameter methods that may excel in specific settings but fail in others, BOOST dynamically adapts its strategy based on observed data characteristics, eliminating the need for manual tuning while ensuring consistent performance across diverse optimization scenarios. Our results demonstrate that adaptive hyperparameter selection can achieve notable performance gains compared to fixed hyperparameter methods.
Appendix A Search Space Description
In this section, we describe the four discretized synthetic benchmark functions and the HPO-B dataset used in our experiments.
Discretized Synthetic Benchmark Functions
To reflect realistic discrete experimental conditions while preserving the core characteristics of each function, we discretized the domain by dividing each axis into 31–41 evenly spaced points. This maintains their optimization complexity while enabling comparison in a discrete setting. The original formulas of the benchmark functions are summarized in Table 2, which serve as the basis for our evaluation.
Function Definition (with $d$: dimension)
Ackley $f(x)=-a\exp\left(-b\sqrt{\frac{1}{d}\sum x_{i}^{2}}\right)-\exp\left(\frac{1}{d}\sum\cos(cx_{i})\right)+a+e$, where $a=20,\ b=0.2,\ c=2\pi$.
Levy $f(x)=\sin^{2}(\pi w_{1})+\sum_{i=1}^{d-1}(w_{i}-1)^{2}[1+10\sin^{2}(\pi w_{i}+1)]+(w_{d}-1)^{2}[1+\sin^{2}(2\pi w_{d})]$,
where $w_{i}=1+\frac{x_{i}-1}{4}$.
Rosenbrock $f(x)=\sum_{i=1}^{d-1}[100(x_{i+1}-x_{i}^{2})^{2}+(1-x_{i})^{2}]$.
SumSquares $f(x)=\sum_{i=1}^{d}i\cdot x_{i}^{2}$.
Table 2: Synthetic Benchmark Functions
  1. 1.
    Ackley Function
    Discretized on $x_{i}\in[-31.5,31.5]$ with 41 points (step size = 1.75) to preserve periodic local minima.
  2. 2.
    Levy Function
    Discretized on $x_{i}\in[-10,10]$ with 31 points to retain misleading humps leading to local minima.
  3. 3.
    Rosenbrock Function
    Discretized on $x_{i}\in[-5,10]$ with 31 points to preserve the narrow valley structure.
  4. 4.
    SumSquares Function
    Discretized on $x_{i}\in[-10,10]$ with 31 points to retain the global bowl-shaped curvature.
Machine Learning Hyperparameter Optimization Tasks
The HPO-B dataset (Arango et al. 2021) is a benchmark suite designed for hyperparameter optimization of machine learning (ML) models. It provides pre-evaluated performance metrics over well-defined hyperparameter search spaces for various ML algorithms across different datasets and dimensionalities. For our experiments, we selected five representative tasks covering a range of dimensions and model types:
  • Decision Trees (6D)
    Search Space IDs: 5636, 5859
    The search space includes six variables:
    • minsplit: Minimum number of observations required to attempt a split
    • minsplit.na: NA indicator for minsplit
    • minbucket: Minimum number of observations in a terminal node
    • cp: Complexity parameter for pruning
    • maxdepth: Maximum depth of any node in the tree
    • maxdepth.na: NA indicator for maxdepth
  • Random Forests (9D)
    Search Space IDs: 7607, 7609
    The search space includes nine variables:
    • num.trees: Number of trees to grow
    • num.trees.na: NA indicator for num.trees
    • mtry: Number of variables sampled at each split
    • min.node.size: Minimum number of observations in a terminal node
    • sample.fraction: Fraction of observations used for each tree
    • respect.unordered.factors.ohe.na: NA indicator for categorical encoding
    • respect.unordered.factors.ohe.TRUE: One-hot encoding indicator
    • replace.ohe.FALSE: Sampling without replacement indicator
    • replace.ohe.na: NA indicator for replacement option
  • XGBoost (16D)
    Search Space ID: 5906
    The search space includes sixteen variables:
    • eta: Learning rate
    • max_depth: Maximum depth of a tree
    • max_depth.na: NA indicator for max_depth
    • min_child_weight: Minimum sum of instance weight in a child
    • min_child_weight.na: NA indicator
    • subsample: Subsample ratio of training instances
    • colsample_bytree: Subsample ratio of columns per tree
    • colsample_bytree.na: NA indicator
    • colsample_bylevel: Subsample ratio of columns per level
    • colsample_bylevel.na: NA indicator
    • lambda: L2 regularization term
    • alpha: L1 regularization term
    • nrounds: Number of boosting rounds
    • nrounds.na: NA indicator
    • booster.ohe.na: NA indicator for booster encoding
    • booster.ohe.gblinear: One-hot encoding indicator for gblinear booster
All data from HPO-B are min–max normalized before evaluation. Input parameters $x$ are rounded to five decimal places (i.e., to the nearest $10^{-5}$). If multiple data points share the same $x$ after rounding, their corresponding $y$-values are averaged to obtain a single value.
Appendix B Information for Reproducibility
All experiments were repeated 10 times to ensure statistical reliability. To ensure reproducibility, we controlled randomness by adopting a consistent seeding strategy: for the $n$th trial, we used seed $n-1$ (i.e., seeds 0–9). For K-means clustering, we fixed the random seed to 42 for consistency with common practice
Hyperparameter Settings
All key hyperparameters used in our experiments are listed below. Additional configuration details, such as the number of initial samples and optimization iterations, are provided in the main text.
  • GP Constraints:
    • Noise constraint: $[5\times 10^{-4},0.2]$
    • Length scale constraint: $[5\times 10^{-6},\sqrt{d}]$, where $d$ is the input dimension
    • Output scale constraint: [0.05,20.0][0.05,20.0][ 0.05 , 20.0 ]
  • GP Optimizer:
    • Optimizer: Adam
    • Learning rate: 0.05
    • Maximum training iterations: 50
Hardware
All experiments were conducted on a CPU under WSL2 (Windows Subsystem for Linux 2), which provides a Linux-compatible environment within the Windows operating system. The hardware used includes an AMD Ryzen 9 7900X CPU and an NVIDIA GeForce RTX 5060 Ti GPU (although all computations in this study were executed on the CPU), with 64 GB of RAM.
Software
The software environment is detailed as follows:
  • Operating System: Ubuntu 20.04.6 LTS (via WSL2)
  • Host OS: Windows 11
  • Python: 3.11.11
  • PyTorch: 2.7.1+cu128 (executed on CPU)
  • torchvision: 0.22.1+cu128
  • torchaudio: 2.7.1+cu128
  • GPyTorch: 1.14
  • BoTorch: 0.14.0
  • NumPy: 2.1.2
  • SciPy: 1.15.3
  • scikit-learn: 1.6.1
  • pandas: 2.2.3
Appendix C Additional results
In the Method section, we proposed several design strategies to enhance the robustness of BOOST:
  1. 1.
    Using K-means clustering during the data partitioning step,
  2. 2.
    Using a $|r_{n}|:|q_{n}|$ ratio of 1:2,
  3. 3.
    Using the target value $y_{target}$ for stopping as the top 5th percentile of function values in $D_{n}$,
  4. 4.
    Applying a rule-based tie-breaking mechanism in the recommendation step.
The following results validate the effectiveness of these design choices.
Effect of Data Partitioning Method
Figure 3 compares two data partitioning strategies: K-means clustering (default) and random selection of the reference subset $r_{n}$ from the data-in-hand $D_{n}$. BOOST with K-means clustering consistently outperformed the variant using random selection, highlighting the importance of structured data partitioning.
Figure 3: Comparison of BOOST performance under different data partitioning strategies
Effect of Data Partitioning Size
Figure 4 compares the optimization performance of BOOST under various $|r_{n}|:|q_{n}|$ ratios, where $q_{n}$ denotes the query subset. No significant performance differences were observed across most ratios, and no particular setting consistently outperformed others.
However, we found that using an excessively small reference set—such as a 1:4 ratio—led to a notable degradation in performance. This is likely due to the increased uncertainty of the internal Gaussian Process model, which hinders fair and stable evaluation during the optimization process.
Figure 4: Optimization performance of BOOST with different $|r_{n}|:|q_{n}|$ ratios.
Effect of Stopping Criteria
Figure 5 compares two stopping criteria used in the internal BO process of BOOST: setting the target value to the top 5th percentile of $D_{n}$ versus using the known global optimum. Using the top 5th percentile consistently outperformed the global optimum criterion. The reason is that when the global optimum is used as the target, a kernel–acquisition pair that performs well overall may still be penalized if it fails to reach the global optimum due to stochasticity. This can lead to an underestimation of its quality and hinder proper selection.
Figure 5: Optimization performance of BOOST with different stopping criteria.
Effect of Tie-Breaking Rule
Figure 6 compares two tie-breaking strategies used in the recommendation step: a predefined rule (see Methods section in the main text) and random selection. The results show that the predefined priority rule consistently yields better performance. This demonstrates that rule-based tie-breaking significantly impacts the outcome by acquiring more informative samples, which in turn helps to reveal the structural characteristics of the objective function.
Figure 6: Optimization performance of BOOST with different tie-breaking rules.